1.Suppose that a body is accelerated at a rate of 2 m/s2.If the net force is tripled and mass doubled, calculate the new acceleration of the body .
2.A car negotiating a curved path , travels one quarter of the circular path of radius ‘r’. Find the distance traveled and the displacement .
3.A car travels a certain distance with a speed of 60 km/h and then returns back to the starting point with a speed of 50 km/h . Find the average speed and average velocity of the body .
1
Expert's answer
2012-07-17T07:47:56-0400
1.
F=ma
3F=2m a1 =>a1=3F/2m=1.5a=1.5*2=3 m/s^2

2. Distance s=2pi r/4=pi r/2
Displacement d=r sqrt(2)

3. Average speed=total distance/time
Let this certain distance be d. Then time traveled with speed 60 is d/60, time traveled back is d/50
So we have average speed=2d/(d/60+d/50)=2d*3000/(110d)=600/11=54,5 km/h
Average velocity=total displacement/time
Since total displacement=0 then average velocity=0